Grade 7Physical Chemistry

What mass of (NH4)2CO3 will produce 3 mole of sufficient acid ? . Pls give the chemical reaction and tell how to identify from given salts which acid will be produced ?

Pushkar Aditya
12 Years agoGrade 7

1 Answer

Sumit Majumdar
12 Years ago
Dear student,
The reaction goes as follows:
[\left ( NH_{4} \right )_{2}CO_{3}\rightarrow 2NH_{3}+H_{2}O+CO_{2}]
The combination of water and carbon di oxide gives carbonic acid.
Hence for producing 3 moles of carbonic acid, we need 3 moles of [\left ( NH_{4} \right )_{2}CO_{3}]
The molar mass of [\left ( NH_{4} \right )_{2}CO_{3}] is equal to:
[\left (14+1\times 4 \right )\times 2+12+16\times 3=36+12+48=96 g/mol]
Hence, the mass required would be equal to:
96 [\times] 3=288 g
Regards
Sumit